OpenStudy (anonymous):

What's the vertex of y=x^2+6x+8?(

OpenStudy (anonymous):

x-coordinate of vertex when y=ax^2+bx+c = -b/2a

OpenStudy (abb0t):

Start by completing the square first.

OpenStudy (anonymous):

Once you have the x-coordinate you can easily solve for y-coordinate by plugging in

OpenStudy (anonymous):

the answer is {-3,-1} have a nice day=D

OpenStudy (abb0t):

HINT: Remember that the vertex form is: \[y = a(x-h)^2+k\]

OpenStudy (anonymous):

aw thanks=D Im really good with problems that don't involve words lol

OpenStudy (abb0t):

giving away the answer doesn't help the person learn, @Helly :P

OpenStudy (anonymous):

SORRY=(

OpenStudy (anonymous):

How do you figure out if its the maximum or minimum?

OpenStudy (anonymous):

Is the curve going up or down? Look at the sign of "a"

OpenStudy (abb0t):

You could take a few derivatives.